# Equitability and explosive synchronisation in multiplex and higher-order networks

### v1

@@ Abstract
Cluster synchronisation is a key phenomenon observed in networks of coupled dynamical units. Its presence has been linked to symmetry and, more generally, to equability of the underlying network pattern of interactions between identical dynamical units. In this article, we clarify once and for all the relation between equitability and cluster synchronisation on a very general dynamical system which allows multi-layer and higher-order interactions. Namely, we show that equitability is a necessary, and sufficient, condition for the existence of independent cluster synchronised solutions. As an important consequence, our results explain the ubiquity of explosive synchronisation, as opposed to cluster synchronisation, in multi-layer and higher-order networks: equitability imposes additional constraints that must be simultaneously satisfied on the same set of nodes. Our results have important implications for the design of complex dynamical systems of coupled dynamical units with arbitrary cluster synchronisation patterns and coupling functions.
@@ I Introduction
The study of synchronization in coupled networks of oscillators Boccaletti et al. (2006) serves as a fundamental pillar to understand the dynamics of interconnected systems, holding profound implications across numerous disciplines Rosenblum and Kurths (2003); Boccaletti et al. (2002, 2018). At its core, the existence of a synchronized state represent how individual elements within a network coordinate their behaviours, fostering coherence or collective dynamics. In recent years, traditional results of global synchronizability in standard networks Pecora and Carroll (1998) have been extended to both systems with multiple layers of interactions Boccaletti et al. (2014); Della Rossa et al. (2020); Del Genio et al. (2016); Boccaletti et al. (2018); Del Genio et al. (2022) as well as systems with many-body interactions Battiston et al. (2020); Gambuzza et al. (2021), constituting a new paradigm that allows researchers to better model the complex processes of real-world systems Boccaletti et al. (2023).
@@
A synchronized state may not involve the whole network. In fact, under specific conditions, one can observe one or more subsets of nodes that synchronize their motion independently of the other groups of nodes and of the rest of the network, giving rise to the phenomenon known as cluster synchronization. Notably, even though the nodes within a specific cluster evolve in a coordinated fashion, displaying similar patterns of activity or dynamics, nodes in other clusters may have different synchronization patterns or even remain desynchronized from each other. Understanding how clusters within a network synchronize their activities impacts multiple fields, as it offers insights into the general processes of organization and dynamics in complex systems Pecora et al. (2014); Schaub et al. (2016); Sorrentino et al. (2016); Siddique et al. (2018); Sánchez-García (2020); Jafari et al. (2024); Sun et al. (2025); Parastesh et al. (2025), including brain networks, social networks, power grids, and biological systems.
Recently, the study of synchronisation has been considered in generalisations of network models, such as multilayer and hypergraph models Del Genio et al. (2016); Gambuzza et al. (2021). In contrast to network models, explosive synchronisation has been commonly observed in higher-order network dynamics Battiston et al. (2021); Ghorbanchian et al. (2021). In this article, we formally explain the role of equitability in cluster synchronisation for a very general system of coupled identical dynamical units in multilayer and hypergraph networks.
In particular, we show that the occurrence of equitability, and thus (generic, or independent) cluster synchronisation, is much more rare in multilayer and higher-order networks, as it must be simultaneously and satisfied on each layer, or at each order of multi-body interaction, independently. This explains why cluster synchronisation is uncommon in such systems, which, instead, almost universally feature explosive synchronisation.
More specifically, we show that equitability (of the partition of the underlying interaction multiplex, or hypergraph, into clusters) is a necessary condition for the existence of independent cluster-synchronised solutions, that is, synchronised solutions for which the coupling functions between clusters are linearly independent over trajectories. Conversely, equitability is also a sufficient condition for synchronisation: we can construct cluster synchronised solutions with respect to any given partition, as long as it is equitable.
@@
The results are independent of the specific form of the coupling functions or of the coupling strength parameters, and depend only of the structural properties (equitability) of the underlying interaction patterns (a network, multiplex, or hypergraph), directly relating structure to function. The caveat is that we only concern ourselves with the existence and not the stability of the synchronised solutions, which, indeed, typically depends on the specific form of the coupling functions and, for fixed coupling functions, on the values of the coupling strength parameters. Nevertheless, our results show how equitability gives strong restrictions on the existence of synchronised solutions, and, in particular, how the restrictions are much harder to achieve for higher-order dynamical interactions.
The existence of cluster synchronised solutions has been previously related to the presence of network symmetries and, more generally, of equitable partitions Schaub et al. (2016) and balanced equivalent relations Golubitsky and Stewart (2006). Our structural equitability results can be seen as an instantiation of the formalism in Golubitsky and Stewart (2006) to multilayer networks, and an extension to many-body interactions (hypergraphs). In addition, we characterise non-equitable solutions in terms of the linear independence of the coupling functions over trajectories, study difference notions of dynamical and structural equitability (fully in the SI), and explain the role of higher-order equitability in explosive synchronisation.
II Dynamical system and cluster synchronisation
@@
where $\mathbf{x}_{i}(t)$ is a $D$-dimensional vector describing the state of the $i$th unit at time $t\in T\subseteq\mathbb{R}$, $\mathbf{f}$ is a smooth map that encodes the internal dynamics of the units, and $\mathbf{h}_{i}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$ is a coupling term, representing the total dynamical input of all other nodes to node $i$. In a multilayer network (strictly speaking, a multiplex network Boccaletti et al. (2014), i.e. same $N$ nodes on each layer) with $M$ layers, the coupling is of the general form
@@
whereas in a higher-order network (hypergraph) with highest multi-body interactions order $M$, it is Gambuzza et al. (2021)
@@
In the equations above, $\sigma_{m}>0$ is the coupling strength for layer $m$ or order $m$, $A^{(m)}$ is the adjacency matrix for layer $m$, or the hypergraph adjacency tensor of order $m+1$, and the coupling functions $\mathbf{g}$ are assumed to be synchronization non-invasive, that is, $\mathbf{g}^{(m)}(\mathbf{x},\mathbf{x},\dotsc,\mathbf{x})=0$ for all $m$ and any $\mathbf{x}$. For $M=1$, both the multilayer and the hypergraph case reduce to the network (or monolayer) case,
@@
A solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ to the dynamical system Eq. (1) is globally synchronised if $\mathbf{x}_{1}=\ldots=\mathbf{x}_{N}$ (identical synchronisation). Since $\mathbf{g}$ is non-invasive, this is equivalent to the synchronised solution $\mathbf{x}_{s}=\mathbf{x}_{i}$ being a solution of $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. Many systems, however, exhibit partial synchronisation, that is, identical synchronisation in a subset, or several subsets, of nodes or dynamical units Pecora et al. (2014); Boccaletti et al. (2018, 2002). Formally, given a solution $\vec{\mathbf{x}}$, we say that a subset of nodes $C\subseteq V$ is synchronised if $\mathbf{x}_{i}=\mathbf{x}_{j}$ for all $i,j\in C$. If $\mathcal{P}$ is a partition of the node set $V$ (a collection of pairwise disjoint subsets of $V$, called clusters, whose union is $V$), we say that a solution $\mathbf{x}$ is $\mathcal{P}$-synchronised if every $C\in\mathcal{P}$ is synchronised, that is,
@@
for each $C\in\mathcal{P}$. This includes any form of partial synchronisation, as we can put single nodes on (trivially) synchronised clusters on their own — we call such cluster a singleton.
@@
For simplicity, we focus on the multilayer case first, and discuss the hypergraph case later, in the Section called ‘Results for hypergraph dynamics’.
III Dynamical and structural equitability
We want to relate cluster (or partial) synchronisation to equitability, that is, the fact that every node receives the same input from nodes in other clusters. We first introduce a notion of dynamical equitability, then relate this to a more standard notion of structural equitability, normally referred to as ‘external equitability’ in the cluster synchronisation literature Schaub et al. (2016); O’Clery et al. (2013).
Given a subset of nodes $C\subseteq V$ and a solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ of the dynamical system, we write
$\mathbf{h}_{i}^{C}$ $=\sum_{m=1}^{M}\sigma_{m}\sum_{j\in C}^{M}A_{i,j}^{(m)}\mathbf{g}^{(m)}(\mathbf{x}_{i},\mathbf{x}_{j})$ (6)
@@ III Dynamical and structural equitability
We can now rewrite Eq. (1) in the multilayer case in terms of the internal and external dynamical inputs with respect to a partition $\mathcal{P}$,
@@
We would like to relate this notion to a more standard notion of structural equitability, that is, equitability with respect to the interaction pattern of the nodes in the underlying multilayer network. If $C\subseteq V$, we write
$h_{i}^{C,m}=\sum_{k\in C}A^{(m)}_{ik}.$ (9)
for the structural input to node $i$ from all nodes in $C$ on layer $m$. This is a constant that depends only on the partition $\mathcal{P}$ and the underlying interaction multilayer network, but not on the coupling functions $\mathbf{g}^{(m)}$ nor on a solution $\mathbf{x}$. We say that $\mathcal{P}$ is structurally equitable if, for all $C,C^{\prime}\in\mathcal{P}$, $C\neq C^{\prime}$, and $1\leq m\leq M$,
@@
for all $i,j\in C$. The condition $C\neq C^{\prime}$ results in ‘external’ equitability, and is included since the internal connectivity in each cluster does not play a role in cluster synchronisation, as shown above. A partition $\mathcal{P}$ satisfying Eq. (10) is often called external equitable partition in the literature, and has been extensively related to cluster synchronisation Schaub et al. (2016); Sorrentino et al. (2016); Golubitsky and Stewart (2006).
It is easy to show that structural equitability implies dynamical equitability for any cluster synchronised solution (see SI). We will show that, under a natural cluster linear independence condition (next section), dynamical and structural equitability are in fact equivalent for cluster-synchronised solutions, and relate them more generally to cluster synchronisation.
In the SI, we study several definitions of dynamical and structural equitability and their relationships in more detail, and justify our choices above.
IV Independent cluster synchronisation
We call a cluster independent if the coupling functions (across clusters, and layers) with all other clusters are linearly independent over trajectories. Formally, given $\mathcal{P}$ a partition and $\vec{\mathbf{x}}$ a $\mathcal{P}$-synchronised solution, we call the solution $\vec{\mathbf{x}}$ $\mathcal{P}$-independent if, for each $C\in\mathcal{P}$, the $M\cdot(|\mathcal{P}|-1)$ functions
@@ IV Independent cluster synchronisation
are linearly independent over the time domain $T$, where we write $\mathbf{x}_{C}(t)$ for the synchronised solution on $C$ (i.e. $\mathbf{x}_{C}(t)=\mathbf{x}_{i}(t)$ for all $i\in C$), and similarly for $\mathbf{x}_{C^{\prime}}$ (see the SI for full details). We call this (linearly) independent cluster synchronisation: cluster synchronisation (Eq. (5)) with linearly independent coupling functions over trajectories (Eq. (11)).
In some sense, independent cluster synchronisation is ‘generic’: clusters are independent unless they are synchronised ($\mathbf{x}_{C}=\mathbf{x}_{C^{\prime}}$, which we can resolve by merging synchronised clusters together), or there is a non-trivial linear relation between them, as we illustrate next.
There are several ways in which a cluster may fail to be independent:
@@
and any combination of the above.
All but case (1), which forces synchronised nodes to be in the same cluster, depend on the specific form of the $M$ coupling functions $\mathbf{g}^{(m)}$, and in particular on their algebraic properties, and thus we cannot say anything more in general. From now on, therefore, we will focus on independent cluster synchronisation.
V Independent cluster synchronisation implies equitability
Our first result is that, under independent synchronisation, dynamical and structural equitability are equivalent. That is, if $\vec{\mathbf{x}}$ is a $\mathcal{P}$-synchronised (Eq. (5)), $\mathcal{P}$-independent (Eq. (11)) solution, then the following are equivalent:

### v3

@@ Abstract
Cluster synchronisation is a key phenomenon observed in networks of coupled dynamical units. Its presence has been linked to symmetry and, more generally, to equability of the underlying pattern of interactions between dynamical units. However, it is not known under which conditions equitability-induced synchronisation is the only cluster synchronisation that can occur on a particular system. Here, we reveal a natural linear independent condition such that equitability becomes necessary, and sufficient, for the existence of cluster synchronised solutions on a very general dynamical system which allows multiplex or higher-order, arbitrarily weighted interactions. Our results explain the ubiquity of explosive synchronisation, as opposed to cluster synchronisation, in multiplex and higher-order networks: equitability imposes additional constraints that must be simultaneously satisfied on the same set of nodes. Our results have significant implications for the design of complex dynamical systems of coupled dynamical units with arbitrary cluster synchronisation patterns and coupling functions.
@@ I Introduction
The study of synchronization in coupled networks of oscillators boccaletti2006structure serves as a fundamental pillar to understand the dynamics of interconnected systems, holding profound implications across numerous disciplines rosenblum2003synchronization ; boccaletti2002synchronization ; boccaletti2018synchronization . At its core, the existence of a synchronized state represent how individual elements within a network coordinate their behaviours, fostering coherence or collective dynamics. In recent years, traditional results of global synchronizability in standard networks pecora1998master have been extended to both systems with multiple layers of interactions boccaletti2014structure ; della2020symmetries ; delgenio2016synchronization ; boccaletti2018synchronization ; delgenio2022mean as well as systems with many-body interactions battiston2020networks ; gambuzza2021stability , constituting a new paradigm that allows researchers to better model the complex processes of real-world systems boccaletti2023structure .
@@
A synchronized state may not involve the whole network. In fact, under specific conditions, one can observe one or more subsets of nodes that synchronize their motion independently of the other groups of nodes and of the rest of the network, giving rise to the phenomenon known as cluster synchronization. Notably, even though the nodes within a specific cluster evolve in a coordinated fashion, displaying similar patterns of activity or dynamics, nodes in other clusters may have different synchronization patterns or even remain desynchronized from each other. Understanding how clusters within a network synchronize their activities impacts multiple fields, as it offers insights into the general processes of organization and dynamics in complex systems pecora2014cluster ; schaub2016graph ; sorrentino2016complete ; siddique2018symmetry ; sanchez2020exploiting ; jafari2024periodic ; sun2025chimera ; parastesh2025synchronization , including brain networks, social networks, power grids, and biological systems. The existence of cluster synchronised solutions has been related to the presence of network symmetries and, more generally, of equitable partitions schaub2016graph ; o2013observability , balanced equivalent relations golubitsky2023dynamics ; golubitsky2005patterns ; stewart2003symmetry ; aguiar2023network and graph fibrations stewart2024dynamics ; boldi2002fibrations .
Recently, the study of synchronisation has been considered in generalisations of network models, such as multilayer and hypergraph models delgenio2016synchronization ; gambuzza2021stability . Indeed, there is increasing evidence of the importance of higher-order interactions in real-world systems battiston2020networks ; delabays2025hypergraph , with emerging applications of higher-order dynamics to real-world systems momani2024directed . In contrast to network models, explosive synchronisation has been commonly observed in higher-order network dynamics battiston2021physics ; ghorbanchian2021higher .
@@
In this article, we formally explain the role of equitability in cluster synchronisation for a very general system of coupled identical dynamical units in multiplex networks (a type of multilayer network with the same nodes on each layer) and hypergraphs (see Fig. 1). We explain why equitability (and thus cluster synchronisation) is much less common in multiplex or higher-order (hypergraph) interactions compared to pairwise (network) interactions. Namely, we show that equitability (of the partition of the underlying interaction multiplex, or hypergraph, into clusters) is a necessary condition for the existence of independent cluster-synchronised solutions (on those clusters), that is, synchronised solutions for which the coupling functions between clusters are linearly independent over trajectories. Conversely, equitability is also a sufficient condition for synchronisation: we can construct cluster synchronised solutions with respect to any given partition, as long as it is equitable, from the quotient dynamics. The results are independent of the specific form of the coupling functions or of the coupling strength parameters, and depend only of the structural properties (equitability) of the underlying interaction patterns (a network, multiplex, or hypergraph). Although we only concern ourselves with the existence and not the stability of the synchronised solutions (which depends on the specific form of the coupling functions and, for fixed coupling functions, on the values of the coupling strength parameters), our results show how equitability gives strong restrictions on the existence of synchronised solutions, and, in particular, how the restrictions are much harder to achieve for higher-order dynamical interactions.
II Results
The results section is organised as follows. After describing the dynamical systems, we first study the multiplex (over the next few sections), then the hypergraph case when the coupling functions are non-invasive, which includes the diffusive case and relates to external equitability. Then we explain how to generalise our results to general coupling functions, which relates to general equitability, and why both cases must be treated separately. We finish with a discussion on explosive synchronisation, and on finding equitable partitions.
II.1 Dynamical system and cluster synchronisation
@@
where $\mathbf{x}_{i}(t)$ is a $D$-dimensional vector describing the state of the $i$th unit at time $t\in T\subseteq\mathbb{R}$, $\mathbf{f}$ is a smooth map that encodes the internal dynamics of the units, and $\mathbf{h}_{i}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$ is a coupling term, representing the total dynamical input of all other nodes to node $i$. In a multiplex network with $M$ layers (and the same $N$ nodes on each layer), the coupling is of the general form
@@
whereas in a higher-order network (hypergraph) with highest multi-body interactions order $M$, it is gambuzza2021stability
@@
(See Fig. 1 for some toy examples.) In the equations above, $\sigma_{m}>0$ is the coupling strength for layer $m$ or order $m$, $A^{(m)}$ is the adjacency matrix for layer $m$ (in the multiplex case), or the hypergraph adjacency tensor of order (number of elements in a relation, in the hypergraph case) $m+1$, and the coupling functions $\mathbf{g}$ are assumed to be synchronization non-invasive, that is, $\mathbf{g}^{(m)}(\mathbf{x},\mathbf{x},\dotsc,\mathbf{x})=0$ for all $m$ and any $\mathbf{x}$. Non-invasiveness jungling2015synchronization ; gambuzza2021stability guarantees that the coupling does not alter the intrinsic dynamics of the dynamical units when they are synchronised, and it is a mild generalisation of the class of diffusive systems, that is, those in which coupling functions depend (possibly linearly) on differences between states. Our results, however, can be adapted to generic coupling functions beyond non-invasive ones; this needs to be treated as a separate case, explained in the section ‘General coupling’ after discussing the non-invasive multiplex, and hypergraph, cases. For $M=1$, both the multiplex and the hypergraph case reduce to the network (or monolayer) case,
@@
A solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ to the dynamical system Eq. (1) is globally synchronised if $\mathbf{x}_{1}=\ldots=\mathbf{x}_{N}$ (identical synchronisation). Since $\mathbf{g}$ is non-invasive, this is equivalent to the synchronised solution $\mathbf{x}_{s}=\mathbf{x}_{i}$ being a solution of $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. Many systems, however, exhibit partial synchronisation, that is, identical synchronisation in a subset, or several subsets, of nodes or dynamical units pecora2014cluster ; boccaletti2018synchronization ; boccaletti2002synchronization . Formally, given a partition $\mathcal{P}$ of the node set $V$ (a collection of pairwise disjoint subsets of $V$, called clusters, whose union is $V$), we say that a solution $\vec{\mathbf{x}}$ is $\mathcal{P}$-synchronised if
@@
for each $C\in\mathcal{P}$. This includes any form of synchronisation (that is, involving any subsets of nodes), as we can put single nodes on (trivially) synchronised clusters on their own — we call such cluster a singleton. We note that, if the partition has $k$ clusters, the case $k=1$ corresponds to global synchronisation and the case $k=n$ to no synchronisation.
@@
For simplicity, we focus on the multiplex case first, and discuss the hypergraph case later, in the Section called ‘Results for hypergraph dynamics’.
Figure 1: Multiplex and hypergraph dynamical systems. Toy example of $\mathbf{a)}$ a biplex (multilayer network with two layers with three identical nodes), and $\mathbf{b)}$ a hypergraph with pairwise and triadic interactions among three nodes, along with their respective adjacency matrices/tensors and dynamical equations. The shorthand $(123)$ indicates any permutation of the three indices. To illustrate equitability, consider the partition into clusters $\{1,3\}$ and $\{2\}$: it is externally equitable for the multiplex $\mathbf{(a)}$ if $w_{12}^{(1)}=w_{23}^{(1)}$ (layer 1) and $w_{12}^{(2)}=w_{23}^{(2)}$ (layer 2); and for the hypergraph $\mathbf{(b)}$ if $w_{12}=w_{23}$.
II.2 Dynamical and external equitability
We want to relate cluster (or partial) synchronisation to equitability, that is, the fact that every node receives the same input from nodes in other clusters. We first introduce a notion of dynamical equitability, then relate this to a more standard notion of external equitability schaub2016graph ; sorrentino2016complete ; o2013observability .
First, we introduce some notation. Given a subset of nodes $C\subseteq V$ and a solution $\vec{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ of the dynamical system, we write
$\mathbf{h}_{i}^{C}$ $=\sum_{m=1}^{M}\sigma_{m}\sum_{j\in C}A_{i,j}^{(m)}\mathbf{g}^{(m)}(\mathbf{x}_{i},\mathbf{x}_{j}),$ (6)
@@ II.2 Dynamical and external equitability
We can now rewrite Eq. (1) in the multiplex case in terms of the internal and external dynamical inputs with respect to a partition $\mathcal{P}$,
@@
We would like to relate this notion to a more standard notion of external equitability, that is, equitability with respect to the interaction pattern of the nodes in the underlying multiplex network. If $C\subseteq V$, we write
$h_{i}^{C,m}=\sum_{k\in C}A^{(m)}_{ik}$ (9)
for the structural input to node $i$ from all nodes in $C$ on layer $m$. This is a constant that depends only on the partition $\mathcal{P}$ and the underlying interaction multiplex network, but not on the coupling functions $\mathbf{g}^{(m)}$ nor on a solution $\mathbf{x}$. We say that $\mathcal{P}$ is externally equitable if, for all $^{\prime}\in\mathcal{P}, CCC\neq C^{\prime}$, and $1\leq m\leq M$,
@@
for all $i,j\in C$. This means that the total, or aggregated, input to a node in $C$ from all nodes in $C^{\prime}$ is the same (see Fig. 1). The condition $C\neq C^{\prime}$ results in ‘external’ equitability, and is included since the internal connectivity in each cluster does not play a role in cluster synchronisation, as shown above, due to the non-invasiveness of the coupling function (which implies $\mathbf{h}_{i}^{\text{int}}=\mathbf{0}$ for all $i$). Note that condition (10) is always satisfied when $\mathcal{P}$ is the partition into one cluster (the condition becomes empty and thus true by convention) or $N$ clusters.
It is easy to show that external equitability implies dynamical equitability for any cluster synchronised solution (see SI). We will show that, under a natural cluster linear independence condition (next section), dynamical and external equitability are in fact equivalent for cluster-synchronised solutions, and relate them more generally to cluster synchronisation (in the two sections after that).
In the SI, we study several definitions of dynamical and external equitability and their relationships in more detail, and justify our choices above.
II.3 Independent cluster synchronisation
We call a cluster-synchronised solution independent if the coupling functions (across clusters, and layers) with all other clusters are linearly independent over trajectories. Formally, given $\mathcal{P}$ a partition and $\vec{\mathbf{x}}$ a $\mathcal{P}$-synchronised solution, we call the solution $\vec{\mathbf{x}}$ $\mathcal{P}$-independent if, for each $C\in\mathcal{P}$, the $M\cdot(|\mathcal{P}|-1)$ functions
@@ II.3 Independent cluster synchronisation
are linearly independent over the time domain $T$, where we write $\mathbf{x}_{C}(t)$ for the synchronised solution on $C$ (i.e. $\mathbf{x}_{C}(t)=\mathbf{x}_{i}(t)$ for all $i\in C$), and similarly for $\mathbf{x}_{C^{\prime}}$ (see the SI for full details). We call this (linearly) independent cluster synchronisation: cluster synchronisation (Eq. (5)) with linearly independent coupling functions over trajectories (Eq. (11)). We note that Eq. (11) is always satisfied if $\mathcal{P}$ is the partition into one cluster (as empty conditions are true by convention) or $N$ clusters (singleton partition), corresponding to global respectively no synchronisation.
In some sense, independent cluster synchronisation is ‘generic’: clusters are independent unless they are synchronised ($\mathbf{x}_{C}=\mathbf{x}_{C^{\prime}}$, which we can resolve by merging synchronised clusters together), or there is a linear relation between them. Namely, the following are the different ways in which a solution may fail to be independent with respect to a partition:
@@
and any combination of the above. Later, we will see that these are the only cases in which cluster synchronisation (and thus dynamical equitability) can occur without requiring external equitability of the underlying partition.
All but case (1), which simply means that synchronised nodes need to be in the same cluster, depend on the specific form of the $M$ coupling functions $\mathbf{g}^{(m)}$, and in particular on their algebraic properties, and thus we cannot say anything more in general. From now on, therefore, we will focus on independent cluster synchronisation.
II.4 Independent cluster synchronisation implies external equitability
Our first result is that, under independent synchronisation, dynamical and external equitability (see previous section) are equivalent. That is, if $\vec{\mathbf{x}}$ is a $\mathcal{P}$-synchronised (Eq. (5)), $\mathcal{P}$-independent (Eq. (11)) solution, then the following are equivalent: